Project

General

Profile

Monday Operations Meeting : 30/3/2015

Agenda:

  • Action Items from previous meeting
    • Check with Andy on the safety implications of the change of hall configuration during search (Steve)
    • minor changes to checklists (Steve) (stands)
    • Enhance Elog DAQ error messages so that physics shifter has a better idea of what went wrong and whether it is serious. Include error message in Elog entry (All)
    • Think about adding an "Analysable" flag to the run conditions spreadsheet for the near term (Chris)
    • Chase the OnMon changes in the MRD action items (Steve)
    • General comment : Please be methodical when going through checklists and ensure that all steps are carried out. (All)
  • Plans for April 11/12 running
    • Data taking plans
    • Training plans
    • Shifters
  • Friday training
  • Longer term plans
    • 11/12Apr15, 18/19Apr15 and 25/26Apr15: DS, CKOV,TOF0,TOF1, EMR, KL, TOF2 (& tracker?)

      *Check the TOF calibration. It will be important to take some data to be sure that the calibration is stable;
      *Proton absorber study
      *First pass of MICE Muon Beam (MMB) commissioning:
      *Start with the currents used in Step I to check the degree of reproducibility of particle rates;
      *Based on the revised simulation of the beam line (Jaroslaw Pasternak), start the search for the optimum magnet settings;
      *Commissioning and calibration of TOF2, KL and EMR will begin as soon as the south-side PRY is complete and the magnets and downstream
      detectors are in their final Step IV positions.

  • AOB
  • Next meeting : 3pm Tuesday 7th April

Phone access : ReadyTalk number

US 866-740-1260

UK 0800-496-0576

Swiss 0800-70-53-51

Access Code: 7368551

Monday Operations Meeting Minutes : 30/3/2015

  • Action Items from previous meeting
    • Check with Andy on the safety implications of the change of hall configuration during search (Steve)
      • Henry pointed out that larger bits of the floor will be moved in April for installation and we should
        discuss the search procedure with Andy
    • minor changes to checklists (Steve) (stands)
  • Review of weekend run

    Yordan reported that the weekend run had a difficult start on Saturday due to misbehaviour of the ISIS beam.
    ISIS was not aware we planned to run, which turned to be due to a miscommunication between Alan Stevens and the
    ISIS shift crew.

    **There is a new procedure for submitting beam requests : a WORD document should be sent, with the MOM electronic
    signature, to Alan to request running. A template for this document exists on the "Operations/MICE Operations Manager" subpage.

    Once we started running data was taken reasonably smoothly. A list of issues encountered will be shown below.

    Sunday running was reasonably smooth.

  • The items from previous MDRs were reviewed
    • Milorad and Pierrick would work on updating the Controls item\
    • Yordan stated that he wouldn't be able to close off MDR1-DAQ-01 yet, but was requested to clarify the item text.
    • Durga would check that MDR2-CMP-01 had been acted upon
    • Paul K would make certain that Rhys was taking care of MDR2-CMP-02 and MDR2-CMP-03
  • The Web whiteboard was reviewed. All items stand.
  • Actions and Lessons from previous running
    • Q123 rack was powered off at the end of the March 11/12 run. This is not what should happen. Should clarify this
      in the checklists (Henry)
    • We should put together a list of signals that we want ISIS to send us so that they can be integrated into the
      Run Control and CDB (Pierrick/Henry)
    • micethin01 had issues and so Run Control was started on micethin02.
    • RATS daemon crashed. Instructions for shifter are to call BLOC. There should be a description for the shifter
      on how to recognise that the RATS daemon had crashed (Henry)
    • CKOV HV settings had not been updated for the June 2014 calibration run. Noted that the old HV settings are still
      in the documentation. Have contacted Lucien to check this and get new documentation (Steve)
    • During a high momentum run, the cooling water temperature went too high and it took too long to notice. There are
      two issues : (1) which variable in the archiver should we be monitoring and (2) checking the alarm took too long
      as the shifters were doing some other things. We decided that alarms take precedence in the control room
    • The rules for leaving the MLCR with only one shifter are not too clear - not helped by the new SHE codes. As a
      working rule we decided that: whether the MLCR is staffed by two or one shifter is up to the judgement of the
      lead shifter if the target is not dipping and no data is being taken. If one shifter leaves the MLCR they should
      take the shifter phone with them. If data is being taken there should always be two shifters in the MLCR.
    • there was an extended discussion about the changeover from the Run Conditions Spreadsheet to relying on the CDB.
      We decided to use the next few April weekend runs to change over, with the intent that we would use the CDB
      only after the last April run. In order for this to happen a few things must be done
      • Scaler and ISIS information needs to go into the CDB (Pierrick)
      • Shifters should monitor the CDB (using the CDBViewer) and compare with the Run Summary Spreadsheet to
        ensure that the same information is contained in both
      • There should be at least one weekend running where the CDB and the spreadsheet show the same information
        without intervention
  • Action Items
    • Check safety of searches in April runs with Andy (Steve)
    • Request to ISIS that we be allowed to dip the target for a reference run on Friday afternoon (Steve)
    • Modify checklists to make sure that Q123 racks are not switched off (Henry)
    • Put together a list of signals that we ISIS to send us to go into the Run Control or CDB (Pierrick/Henry)
    • Document RATS daemon crash signs in beamline documentation (Henry)
    • Work towards replacement of run conditions spreadsheet (All)

6200.xls (20 KB) Bayes, Ryan, 30 March 2015 16:35

6300.xls (9 KB) Bayes, Ryan, 30 March 2015 16:35